Find the inverse of f(x)=-3+ln(x-5).
Answer by rapture(86)

f(x) = -3+ln(x-5)

Let f(x) = y

y = -3+ln(x-5)

Exchange x and y

x = -3+ln(y-5)

Add 3 to both sides.

x + 3 = ln(y - 5)


Raise both sides with e as the base.


e^(x + 3) = e^(ln(y - 5))

e^(x + 3) = y - 5

Add 5 to both sides.

e^(x + 3) + 5 = y

Replace y with the inverse notation.

Answer: f^(-1) x = e^(x + 3) + 5